OpenStudy (anonymous):

Four acids are analyzed and their Ka values are determined. Which of the following values represents the strongest acid?

OpenStudy (anonymous):

what are the choices?

rvc (rvc):

please give full info about question @Olea0513

OpenStudy (anonymous):

Ka = 2.2 × 10−13 Ka = 6.2 × 10−8 Ka = 1.7 × 10−5 Ka = 1.2 × 10−

rvc (rvc):

hmm... thats not so clear @Olea0513

OpenStudy (anonymous):

thats the question and the choices and im so stumped :(

OpenStudy (anonymous):

The one with the lowest numerical value has the highest pH. pH represents the number of hydrogen ions in the given substance. The lower the pH the more acidic the substance. The general idea is that something with a pH of 1 is highly acidic, 2 a little less acidic and so on, until you reach 7 which is neutral (water). Above 7, substances are no longer acidic, but 'basic.' so which one do you think it is..?

rvc (rvc):

strong acids have Ka values greater then one

rvc (rvc):

so values greater than one is a indication of strong acid

OpenStudy (anonymous):

so the value greater then 1.2 & 1.7 or 2.2 is the strongest acid ??...

rvc (rvc):

hmm.. which is the biggest value

OpenStudy (anonymous):

the lowest values are more acidic than the higher ones, so say it was 1.5 and 1.6, 1.5 is more acidic by a little bit

OpenStudy (anonymous):

OH ! SO 1.2 WAS MORE ACIDIC (:

OpenStudy (anonymous):

yep, should be. make sure to close your question when your done.

rvc (rvc):

no

rvc (rvc):

wait those are for pH values @Olea0513 @felipecasal

OpenStudy (anonymous):

yeah isnt that what their going for?

rvc (rvc):

hmm.. Ka is dissociation constant for weak acids